# Information Sifting Funnel: Privacy-preserving Collaborative Inference Against Model Inversion Attacks

### v1

@@ Abstract.
The complexity of neural networks and inference tasks, combined with the demand for computational efficiency and real-time feedback, poses significant challenges for resource-constrained edge devices. Collaborative inference addresses this by delegating shallow feature extraction to edge devices and uploading feature to the cloud for subsequent inference tasks, reducing computational pressure. However, the transmitted features remain vulnerable to model inversion attacks (MIAs), which can reconstruct original input data. Existing defenses, including perturbation and information bottleneck techniques, provide explainable protection but are limited by the lack of standardized criteria for evaluating MIA difficulty, challenges in mutual information estimation, and trade-offs among usability, privacy, and deployability.
To overcome these limitations, we establish the first criterion for assessing the difficulty of MIAs in collaborative inference, supported by theoretical analysis of existing attacks and defenses and validated through experiments using the Mutual Information Neural Estimator (MINE). Building on these findings, we propose SiftFunnel, a privacy-preserving framework for collaborative inference. Specifically, we first train the edge model with both linear and non-linear correlation constraints to limit redundant information in transmitted features, enhancing privacy protection. Moreover, label smoothing and a cloud-based upsampling module are incorporated to balance usability and privacy. To further support deployability, the edge model is designed with a funnel-shaped structure and integrates attention mechanisms, preserving both privacy and usability. Extensive experiments show that SiftFunnel outperforms state-of-the-art defenses against MIAs, achieving superior privacy protection with less than 3% accuracy loss and an optimal trade-off among usability, privacy, and practicality.
@@ 1. Introduction
The rapid advancement of deep learning has led to increasingly complex and powerful neural network models, enabling remarkable performance in tasks such as image detection and recognition, speech processing, and natural language understanding. However, these advancements also pose significant challenges for edge computing devices (Huckelberry et al., 2024). In fields such as intelligent transportation, autonomous driving, facial recognition, IoT applications involving cameras and sensors, and remote healthcare monitoring, edge devices must process data instantly while handling growing computational demands. The limited resources of edge devices make it difficult for them to efficiently perform complex deep learning inference tasks independently (Shlezinger and Bajić, 2022).
@@
Cloud computing enables inference by uploading data to cloud-based models, offering significant computational power. However, it faces two key limitations: bandwidth and network latency hinder its ability to meet real-time application requirements, and uploading raw data raises privacy concerns, especially in sensitive scenarios. Collaborative inference (CI) (Shlezinger et al., 2021) has emerged as a viable solution to these challenges. By partitioning a neural network into $f_{\text{edge}}$ and $f_{\text{cloud}}$, the edge device processes raw inputs into feature representations via $f_{\text{edge}}$, which are then uploaded to the cloud for further inference by $f_{\text{cloud}}$. This approach balances real-time performance and computational efficiency. In recent years, this technique has garnered attention in practical applications, such as unmanned aerial vehicles (UAVs) (Qu et al., 2023) and IoT systems (Shlezinger and Bajić, 2022), as well as in solutions like Apple’s private cloud computing (PCC) (Engineering and Architecture, 2024). With advancements in 5G, IoT, and AI, collaborative inference holds promise for broader adoption and future impact.
However, recent studies have revealed significant privacy risks in collaborative inference, where input data on edge devices can be reconstructed through model inversion attacks (MIAs) (He et al., 2019; Yang et al., 2022; Yin et al., 2023), as shown in Figure 1. MIAs are fine-grained privacy attacks that leverage the output information of AI models and deep learning generation techniques to reconstruct input data (He et al., 2019; Yang et al., 2019), recover representative training data (Zhang et al., 2020; Liu et al., 2024), or expose sensitive information (Fredrikson et al., 2014; Mehnaz et al., 2022). These risks are particularly severe in collaborative inference scenarios. The root cause is that, compared to centralized model deployment, edge devices typically employ shallow neural networks to compute and upload feature representations. These features contain far more information than the final inference results, enabling adversaries to reconstruct the input data (Azizian and Bajić, 2024). For instance, in facial recognition applications, while adversaries in centralized deployments might only recover training data through the model’s API (Yuan et al., 2023; Struppek et al., 2022), those in collaborative inference can directly reconstruct the raw facial data collected by edge devices from the transmitted features.
[…]
Firstly, existing defense techniques in collaborative inference lack a thorough and unified analysis of the fundamental nature of MIA. Current mainstream defense methods can be categorized into three types based on their technical approaches: perturbation-based (He et al., 2020; Wang et al., 2022), information bottleneck (IB)-based (Wang et al., 2024; Peng et al., 2022), and neural network depth & information bottleneck (NND & IB)-based defense strategies (Azizian and Bajić, 2024; Ding et al., 2024). Perturbation-based strategies aim to increase the difficulty of attacks by introducing perturbations to the gradients or output information of the trained model. IB-based strategies focus on approximating the reduction of mutual information $I(x,z)$ between the input data $x$ and the feature $z$, thus minimizing redundant information about $x$ in $z$ (Wang et al., 2021). NND & IB-based strategies leverage the observation that increasing the depth of the edge network can induce an information bottleneck effect, thus enhancing defense. These methods provide theoretically interpretable defense mechanisms, with the introduction of information theory concepts offering notable contributions to defense design. However, existing work lacks a comprehensive understanding of MIA, relies on weak assumptions about adversarial capabilities, and fails to establish a clear criterion for evaluating the complexity of implementing MIA. As a result, defenses against MIAs still have room for improvement, including addressing potential future advancements.
Secondly, the trade-off between usability, privacy, and deployability in edge models remains a challenging and underexplored area in current research. Encryption-based defense strategies receive limited attention in collaborative inference due to resource constraints and the need for multitask processing in real time (Mireshghallah et al., 2021). Additionally, encryption does not guarantee complete data security, as modern side-channel attacks leveraging deep learning can recover encrypted data (Hu et al., 2023). The three main defense strategies discussed above also face issues of imbalance. Perturbation-based methods often compromise usability, whereas IB-based defenses aim to reduce mutual information $I(x,z)$, providing theoretical protection against simpler attacks such as MLE-based MIA. However, accurately computing $I(x,z)$ is challenging, as it relies on approximations through sufficient data and specific algorithms, making it less effective against more advanced attacks such as Gen-based MIA. Adversarial training has been introduced as a component to address the aforementioned issues, but its strong specificity and limited interpretability make it difficult to prevent future advanced attacks while significantly increasing the difficulty of training edge models. The NND & IB-based approach aims to balance privacy and usability by increasing model depth, but this approach significantly raises the resource demands on edge devices, including memory, storage, and bandwidth, thereby reducing deployability. As a result, such methods are often less favored in practical applications.
@@
Thus, there is a pressing need for a novel approach grounded in an in-depth analysis of attack theories to provide interpretable and effective privacy protection while maintaining the usability and deployability of edge models. This need gives rise to three key technical challenges. Challenge 1: How to establish an experimentally validated criterion for assessing MIA difficulty in collaborative inference? Prior studies have interpreted MIA from an information-theoretic perspective, but challenges in computing mutual information and conducting an in-depth analysis of attack principles have hindered the establishment of a verifiable criterion, which is essential for guiding defense design and evaluating methods. Challenge 2: How to effectively and interpretably remove redundant information from transmitted features while preserving edge model usability? Although information bottleneck theory provides an interpretable defensive effect, mutual information is challenging to calculate, and current approximation methods are suboptimal. Challenge 3: How to balance usability and privacy while ensuring practical deployability? NND & IB-based methods significantly increase computational overhead on edge devices, making them challenging for practical application. Moreover, when applied to edge models that perform only simple feature extraction, these methods also struggle to resist advanced Gen-based MIAs.
To overcome the aforementioned challenges, we established a criterion $D_{mia}$ for evaluating the difficulty of MIAs in CI and proposed a lightweight privacy-preserving strategy called Information Sifting Funnel (SiftFunnel), incorporating three key features. (1) SiftFunnel provides interpretable defense principles based on $D_{mia}$ We identified factors influencing attack difficulty in collaborative inference, including $I(x,z)$ and the effective information mean $\delta(z)$, i.e., the average number of non-zero elements in $z$. By analyzing these factors, we clarified defense design directions and developed a framework to address potential future threats. Using the Mutual Information Neural Estimator (MINE) (Belghazi et al., 2018), we quantified the changes in $I(x,z)$ at different network depths, validating the criterion. (2) SiftFunnel effectively reduces redundant information in transmitted features while preserving edge model usability. Guided by the proposed criterion, we enhanced the model’s loss function to include distance-correlation and Pearson-correlation measures, alongside an $l_{1}$-norm constraint on $z$. These refinements reduce $H(z)$, $\delta(z)$ while increasing $H(x|z)$. To further filter redundant information, we integrated Squeeze-and-Excitation (SE) Modules (Hu et al., 2018), Triple Attention Mechanisms (Misra et al., 2021), and Convolutional Block Attention Modules (CBAM) (Woo et al., 2018) into the edge model to focus on task-relevant features. Usability was preserved through a combination of label-smoothing and KL-divergence loss. (3) SiftFunnel balances usability and privacy while light-weighting the edge model. A funnel-shaped edge model was designed to progressively reduce transmitted channels, improving efficiency and reducing redundant information while minimizing memory and storage demands. To maintain usability, an upsampling module was added on the cloud side.
Based on prior evaluation metrics, this paper incorporates $I(x,z)$, quantified using MINE, and $\delta(z)$ as key metrics to validate the theoretical analysis and assess the effectiveness of defense methods. To evaluate deployability, we consider the edge model’s parameter count, memory usage, storage consumption, and inference latency. Through extensive experiments, we demonstrate the validity of the proposed difficulty criterion and the effectiveness of SiftFunnel in balancing usability, deployability, and data privacy compared to existing methods.
@@
  • We establish the first difficulty criterion for implementing MIA in collaborative inference, providing guidance for defense design and preventing future advanced MIAs.
  • We propose SiftFunnel, a privacy-preserving scheme for collaborative inference that constrains redundant transmitted features through linear and nonlinear correlation, ensuring model usability with label smoothing and loss function regularization.
  • SiftFunnel adopts a funnel-shaped structure with integrated attention mechanisms to balance usability and privacy while maintaining the deployability of edge models.
  • We conduct systematic experiments validating the proposed criterion and demonstrating SiftFunnel’s advantages in balancing usability, privacy, and deployability.
@@ 2.1. Collaborative inference
The rapid development of artificial intelligence has enabled deep learning models to achieve exceptional performance in image processing and other tasks. However, the computational demands of these models often exceed the capabilities of resource-constrained edge devices, which face limitations in processing power, storage, and energy efficiency. To address these challenges, collaborative inference (Shlezinger et al., 2021) has emerged as an effective solution. This approach divides the inference process into two stages: one part is performed locally on the edge device (which may consist of a single device or a chain of interconnected devices), while the remaining computation is uploaded to the cloud. Formally, this can be expressed as
@@
Collaborative inference typically begins with training $f$ on centralized datasets in the cloud, followed by partitioning and deploying the $f_{e}$ to the edge. Alternatively, methodologies such as Split Learning  (Zhu et al., 2023) enable edge devices to participate in model training locally. Regardless of the training approach, the inference process remains consistent. As depicted in Figure 1, the edge-deployed model first computes an intermediate representation $feature=f_{e}(x)$, which is then transmitted to the cloud for completion of the inference process $result=f_{c}(feature)$. The final output is either sent back to the edge device or retained in the cloud for further analysis.
Compared to the traditional approach of uploading raw data to the cloud for inference, collaborative inference offers distinct advantages. First, it reduces communication overhead by allowing the transmitted features to be processed, potentially making them smaller in size than the original data. Second, it improves privacy protection by limiting the exposure of sensitive information by transmitting processed features instead of raw data, such as physiological signals or facial images. These advantages make collaborative inference a suitable solution for applications in healthcare, facial recognition, and industrial IoT, where efficiency and privacy are critical requirements.
However, existing research highlights that collaborative inference models are susceptible to Model Inversion Attacks (MIA), where adversaries attempt to reconstruct the original input data from transmitted feature. Notably, in edge-based inference, computations can be distributed across multiple devices in a chained architecture (He et al., 2019), expressed as $f=f_{e1}\circ f_{e2}...\circ f_{c}$. Studies show that deeper neural networks exhibit greater resistance to data reconstruction, making devices positioned closer to the cloud less vulnerable. In contrast, the initial edge device, which typically employs a shallow neural network for feature computation and transmission, generates representations that retain more information from the original input, increasing its susceptibility to MIA. Given this vulnerability, this paper focuses on the MIA threat to the initial edge device, which uses a shallow neural network while the rest resides in the cloud. This common scenario in real-world applications poses a significant security risk due to the heightened likelihood of input data reconstruction, underscoring the need to address this challenge.
@@ 2.2. Model inversion attack
Model Inversion Attack (MIA) is a fine-grained privacy attack that uses the output of a target model and adversarial knowledge, such as white-box access or training data distribution, to reconstruct training data (Zhang et al., 2020), sensitive attributes (Mehnaz et al., 2022), or input data (Yang et al., 2019). Initially proposed by Fredrikson et al. (Fredrikson et al., 2015), MIA was applied to linear regression models, successfully reconstructing genomic data from warfarin dosage predictions and patient attributes, thereby revealing the privacy risks. Subsequent studies (Fredrikson et al., 2015) extended this concept to deep learning, showing that training data from deep neural networks could also be reconstructed.
Current MIA techniques in collaborative inference can be broadly categorized into Maximum Likelihood Estimation-based (MLE-based) MIAs and Generative Model-based (Gen-based) MIAs, which are further analyzed in the following sections.
MLE-based MIAs. MLE-based MIA originated from the 2015 study by Fredrikson et al. (Fredrikson et al., 2015), which reconstructed training data from a multilayer perceptron (MLP). This method minimizes the difference between target and output class confidence scores as a loss function and employs gradient descent to optimize the input, enabling the reconstruction of the original training data. Subsequent research enhanced this approach by incorporating techniques like Total Variation (TV) (Rudin et al., 1992) loss as a regularization term to improve optimization, with these improved methods later referred to as rMLE-based MIA by He et al. (He et al., 2019) In collaborative inference scenarios, MLE-based methods are particularly effective for reconstructing data from shallow neural networks deployed on edge devices. Unlike the prediction vectors produced by full neural networks, the richer feature representations and lower complexity of shallow networks provide adversaries with more detailed and easily optimizable gradient information.
Gen-based MIAs. This method originated from the 2019 work of Yang et al. (Yang et al., 2019), who reconstructed input data in facial recognition systems by using generative deep learning models $g$ to approximate the inverse mapping function of the target model $f^{-1}$, such that $g\approx f^{-1}$. By inputting prediction vectors into $g$, they successfully generated facial input images. Building on this approach, He et al. (He et al., 2019) extended it to collaborative inference scenarios, demonstrating that input data from edge models is more vulnerable to reconstruction, thereby highlighting the risks associated with collaborative inference. Unlike MLE-based methods, Gen-based MIA does not require white-box access to the edge model. Instead, it trains the attack model by calculating losses based on transmitted features, enabling it to approximate the inverse mapping of non-linear computations. Furthermore, as edge models in collaborative inference are typically shallow neural networks, their outputs often retain redundant information from the original input and features, maintaining a strong correlation between inputs and outputs and facilitating the reconstruction process.

### v4

# How Breakable Is Privacy: Probing and Resisting Model Inversion Attacks in Collaborative Inference

@@ Abstract.
Collaborative inference (CI) improves computational efficiency for edge devices by transmitting intermediate features to cloud models. However, this process inevitably exposes feature representations to model inversion attacks (MIAs), enabling unauthorized data reconstruction. Despite extensive research, there is no established criterion for assessing the difficulty of MIA implementation, leaving a fundamental question unanswered: What factors truly and verifiably determine the attack’s success in CI? Moreover, existing defenses lack the theoretical foundation described above, making it challenging to regulate feature information effectively while ensuring privacy and minimizing computational overhead. These shortcomings introduce three key challenges: theoretical gap, methodological limitation, and practical constraint.
To overcome these challenges, we propose the first theoretical criterion to assess MIA difficulty in CI, identifying mutual information, entropy, and effective information volume as key influencing factors. The validity of this criterion is demonstrated by using the mutual information neural estimator. Building on this insight, we propose SiftFunnel, a privacy-preserving framework to resist MIA while maintaining usability. Specifically, we incorporate linear and non-linear correlation constraints alongside label smoothing to suppress redundant information transmission, effectively balancing privacy and usability. To enhance deployability, the edge model adopts a funnel-shaped structure with attention mechanisms, strengthening privacy while reducing computational and storage burdens. Experiments show that, compared to state-of-the-art defense, SiftFunnel increases reconstruction error by $\sim$30%, lowers mutual and effective information metrics by $\geq$50%, and reduces edge burdens by almost 20×20$\times$20 ×, while maintaining comparable usability.
@@ 1. Introduction
The rapid advancement of deep learning has led to increasingly complex neural networks that achieve remarkable performance in tasks like image recognition and object detection. However, deploying such models on edge devices remains challenging due to limited computational resources (Huckelberry et al., 2024). Applications like autonomous driving, facial recognition, IoT surveillance, and remote healthcare require real-time processing, yet the limited computational capacity of edge devices makes standalone deep learning inference impractical (Shlezinger and Bajić, 2022). Cloud computing provides substantial computational power but faces latency, bandwidth, and privacy concerns when transmitting raw data. Collaborative inference (CI) (Shlezinger et al., 2021) mitigates these issues by partitioning neural networks into an edge model $f_{\text{edge}}$ and a cloud model $f_{\text{cloud}}$, where $f_{\text{edge}}$ extracts feature representations for $f_{\text{cloud}}$ to inference. This method enhances efficiency while maintaining real-time performance and is widely applied in UAVs (Qu et al., 2024), IoT systems (Shlezinger and Bajić, 2022), and private cloud computing (PCC) (Engineering and Architecture, 2024). With the rise of 5G, IoT, and AI, CI is poised for broader adoption.
However, recent studies have revealed significant privacy risks in CI, where input data on edge devices can be reconstructed by model inversion attacks (MIAs) (He et al., 2019; Yang et al., 2022; Yin et al., 2023), as shown in Figure 1. Unlike centralized deployments, where adversaries recover training data through model APIs (Yuan et al., 2023; Struppek et al., 2022), CI often relies on shallow neural networks at the edge, producing feature representations that retain significant mutual information (MI) with input data (Azizian and Bajić, 2024). This exposes raw inputs to adversaries, who can reconstruct sensitive data, such as facial images directly from transmitted features (Ding et al., 2024).
MIAs in CI can be categorized into maximum likelihood estimati-on-based (MLE-based) MIAs (He et al., 2019, 2020) and generative model-based (Gen-based) MIAs (Yin et al., 2023; He et al., 2020). MLE-based attacks optimize input data to minimize the loss between the edge model’s output and transmitted features but require white-box access, limiting practicality. Gen-based MIAs, on the other hand, use generative models to approximate the inverse mapping of the edge model, reconstructing inputs from features. Their success depends on data distribution alignment and the model’s ability to learn the inverse function. These privacy risks have driven extensive research on attacks and defenses, but existing methods still face notable limitations.
@@
Firstly, existing research lacks a comprehensive analysis of the fundamental factors influencing MIA effectiveness in CI. Current attack methods focus on improving data reconstruction under specific assumptions, leveraging innovative techniques to enhance reconstruction fidelity and generalization. In recent years, defenses have increasingly recognized the role of mutual information in MIAs and have shifted toward designing optimal information bottleneck (IB) approaches (Wang et al., 2024; Peng et al., 2022) to counter MIAs. However, these methods remain a methodological extension based on existing findings rather than advancing a fundamental theoretical understanding of MIAs. Consequently, their improvements are inherently constrained by existing technical limitations. Identifying the key factors that control MIA effectiveness and developing a systematic criterion for assessing attack difficulty would significantly deepen our theoretical insights and offer broader theoretical guidance for designing more effective defense strategies.
Secondly, the trade-off between usability, privacy, and deployability in edge models is an underexplored and challenging area in current research. Encryption-based defenses receive little attention in CI due to resource constraints, real-time demands (Mireshghallah et al., 2021), and vulnerability to side-channel attacks (Hu et al., 2024). Mainstream MIA defense strategies in CI, constrained by a limited understanding of MIA effectiveness factors, primarily focus on reducing the mutual information $I(x,z)$ between inputs $x$ and extracted features $z$ to counter MIAs (Wang et al., 2024; Peng et al., 2022; Azizian and Bajić, 2024). However, mutual information estimation is inherently challenging due to its reliance on accurate probability density estimation, large-scale training data, and the computational overhead associated with optimization techniques (Wang et al., 2024; Duan et al., 2023), making the usability-privacy trade-off inefficient and difficult to guarantee. Moreover, some approaches compromise usability (He et al., 2020), increase computational and storage burdens on edge devices (Wang et al., 2024), and complicate further fine-tuning and optimization. Consequently, such methods are often less favored in practical applications.
@@
Thus, there is a pressing need for a novel approach grounded in an in-depth analysis of attack theories to provide interpretable and effective privacy protection while maintaining the usability and deployability of edge models. This need gives rise to three challenges. Challenge 1 (Difficulty assessment): How to establish a verifiable criterion for assessing MIA difficulty in CI? The fundamental factors determining MIA effectiveness remain poorly understood, making it difficult to develop a reliable assessment criterion. While MI’s role in MIAs is recognized, its accurate estimation for validation remains challenging. Challenge 2 (Methodological design): How to effectively and interpretably remove redundant information from transmitted features while preserving edge model usability? Although IB theory provides a theoretically sound approach, its practical implementation is limited by the technical constraints of MI estimation. Current methods are computationally expensive and lack flexibility, making fine-tuning and further optimization difficult after initial deployment. Challenge 3 (Practical constraint): How to balance usability and privacy while ensuring practical deployability? Certain defense strategies (Azizian and Bajić, 2024) inevitably increase computational and storage burdens on edge devices. Furthermore, when applied to edge models designed for simple feature extraction, these methods struggle to counter advanced Gen-based MIAs effectively.
To overcome the above challenges, we proposed a criterion $D_{mia}$ to evaluate MIA difficulty in CI and proposed a lightweight privacy-preserving strategy called Information Sifting Funnel (SiftFunnel), featuring 3 key aspects. (1) Criterion basis: SiftFunnel offers interpretable defense principles based on $D_{mia}$ We analyze MIAs using formulaic principles, information theory, and Fano’s inequality, and derive an evaluation formula involving 4 critical factors: mutual information $I(x,z)$, conditional entropy $H(x\mid z)$, feature entropy $H(z)$, and effective information mean $\delta(z)$ (the average number of non-zero elements in $z$). Their impact on MIA performance is validated via the Mutual Information Neural Estimator (MINE) (Belghazi et al., 2018). (2) Usability-privacy trade-off: SiftFunnel effectively reduces redundant information while preserving edge model usability. Aligning with $D_{mia}$ to strengthen privacy protection, we aim to reduce $H(z)$ and $\delta(z)$ while increasing $H(x\mid z)$. Specifically, we incorporate distance correlation, Pearson correlation constraints, and an $l_{1}$-norm constraint on $z$ to enforce these objectives. To maintain usability, we integrate label smoothing and KL-divergence loss into the training process. This optimization enables the edge model to minify redundant information while ensuring the extraction of task-relevant features. (3) Lightweight design: SiftFunnel balances usability and privacy while light-weighting the edge model. We adopt a funnel-shaped edge model to reduce computational and storage overhead. Attention blocks enhance feature selection, while an upsampling module on the cloud side compensates for lost details, preserving inference accuracy. This design, guided by $D_{mia}$, also improves privacy by limiting feature $\delta(z)$ exposure.
Except for previous evaluation metrics, we incorporate MI and $\delta(z)$ as metrics to assess defense effectiveness. Deployability is measured by the edge model’s parameter count and inference latency. Systematic experiments confirm that SiftFunnel effectively balances usability, deployability, and privacy compared to existing methods, achieving superior MIA resistance while reducing mutual information and storage, with an average accuracy trade-off of 3%.
@@
  • We propose the first MIA difficulty criterion in CI, identifying $I(x,z)$, $H(x|z)$, $H(z)$ and $\delta(z)$ as factors impacting attack effectiveness, deepening insights into MIAs.
  • We propose SiftFunnel, a privacy-preserving scheme for CI that constrains redundant transmitted features through linear and nonlinear correlation, ensuring model usability with label smoothing and loss function regularization.
  • SiftFunnel adopts a funnel-shaped structure with attention blocks to balance usability and privacy while ensuring the lightweight design of edge models.
  • We conduct systematic experiments validating the proposed criterion and demonstrating the effectiveness of SiftFunnel in balancing usability, privacy, and deployability.
@@ 2.1. Collaborative inference
Collaborative inference (Shlezinger et al., 2021) has emerged as an effective solution. This approach divides the inference process into two stages: one part is performed locally on the edge device (which may consist of a single device or a chain of interconnected devices), while the remaining computation is uploaded to the cloud. Formally, this can be expressed as
@@
Collaborative inference typically begins with training $f$ on centralized datasets in the cloud, then partitioning and deploying the $f_{e}$ to the edge. Alternatively, approaches like Split Learning  (Zhu et al., 2023) allow edge devices to participate in model training locally. Regardless of the training approach, the inference process remains consistent. As depicted in Figure 1, the edge model computes an intermediate representation, which is transmitted to the cloud for completion of the inference process. The final output is either returned to the edge device or retained in the cloud for further analysis.
However, existing research (He et al., 2019) highlights that edge models are susceptible to MIAs, where adversaries reconstruct input data from transmitted features. Notably, in edge-based inference, computations can be distributed across multiple devices in a chained architecture (He et al., 2019), expressed as $f=f_{e1}\circ f_{e2}...\circ f_{c}$. Studies (He et al., 2020; Ho et al., 2024) show that deeper neural networks closer to the cloud transmit more complex, decision-focused features, making them more resistant to reconstruction. In contrast, the initial edge device generates features retaining more input information, making it highly susceptible to MIAs. This paper focuses on the MIA threat to the initial edge device, which employs a shallow neural network while the remaining computation occurs in the cloud. This common scenario in real-world applications presents a significant security risk due to the increased likelihood of input data reconstruction, highlighting the need for effective countermeasures.
@@ 2.2. Model inversion attack
Current MIA techniques in CI can be broadly categorized into MLE-based MIAs and Gen-based MIAs, with related research progress described as follows.
MLE-based MIAs. MLE-based MIA was first introduced by Fredrikson et al. (Fredrikson et al., 2015), who reconstructed training data from a multilayer perceptron (MLP) by minimizing the difference between target and confidence scores using gradient descent. Subsequent research (He et al., 2019) incorporated Total Variation (TV) (Rudin et al., 1992) loss as a regularization term to improve optimization, with these advanced techniques referred to as rMLE-based MIA. In this paper, both are collectively referred to as MLE-based MIA. In CI scenarios, MLE-based methods are particularly effective for reconstructing data from shallow neural networks, as the richer feature representations and lower complexity provide adversaries with detailed gradient information that is easier to optimize compared to the prediction vectors of deeper networks.
Gen-based MIAs. This approach was introduced by Yang et al. (Yang et al., 2019), who reconstructed input data in facial recognition systems by using generative deep learning models $g$ to approximate the inverse mapping function of the target model $f^{-1}$, such that $g\approx f^{-1}$. By feeding prediction vectors into $g$, they successfully generated facial input images. Building on this approach, He et al. (He et al., 2019) extended it to collaborative inference, demonstrating that input data from edge models is more vulnerable to reconstruction, thereby highlighting the risks. Unlike MLE-based methods, Gen-based MIA does not require white-box access to the edge model. Instead, it trains an attack model by calculating losses based on transmitted features and auxiliary data, enabling it to approximate the inverse mapping of non-linear computations. Furthermore, as edge models produce rich and redundant outputs, this facilitates the training of attack models and the reconstruction process, as analyzed in detail in Section 4.